OpenStudy (anonymous):

Write an equation in slope-intercept form of the line that passes through (– 5, 4) and is parallel to the graph of y = -4x + 1

OpenStudy (jdoe0001):

hmmm what's the slope of \(\bf y = -4x + 1\) anyway?

OpenStudy (anonymous):

Lol this was the exact problem given to me to solve, i dont know

OpenStudy (jdoe0001):

\(\bf \large y = {\color{brown}{ -4}}x + 1\) see the slope now? what do you think it's?

OpenStudy (anonymous):

oh lmao thats what you were asking, i thought you were telling me something was missing

OpenStudy (anonymous):

yes i know what the slope is, and b is 1

OpenStudy (aaronq):

b is the y-intercept. In y=mx+b, the slope is m

OpenStudy (anonymous):

oh so the points are (-5, 4) and (4, 1)?

OpenStudy (aaronq):

nope (4,1) is not a point

OpenStudy (anonymous):

lol if i can just get a full asnwer i can understand it, i am not looking for answers, i am looking to understand

OpenStudy (jdoe0001):

so... .what do you think is the slope?

OpenStudy (anonymous):

ohhhh its 4/1?

OpenStudy (jdoe0001):

well.. notice the negative in front of it \(\bf \bf \large y = {\color{brown}{ -4}}x + 1\)

OpenStudy (anonymous):

oh yeah lol

OpenStudy (anonymous):

-4/1

OpenStudy (jdoe0001):

so a slope parallel to that one will have the same slope as that one so now we know the parallel line will have a slope of -4 and passes through (-5, 4) thus \(\bf \begin{array}{lllll} &x_1&y_1\\ &({\color{red}{ -5}}\quad ,&{\color{blue}{ 4}})\quad \end{array} \\\quad \\ slope = {\color{green}{ m}}= -4 \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (anonymous):

y - 4 = -4 (x + 5) y - 4 = -4x + 5 Add 4 on both sides Y = -4x + 9?

OpenStudy (anonymous):

oh my bad wait

OpenStudy (anonymous):

y - 4 = -4x -20 Add 4 on both sides y = -4x - 16